OpenStudy (anonymous):

Salina purchased a dining room set for $1,730 using a 12-month deferred payment plan with an interest rate of 23.45%. She did not make any payments during the deferment period. What will the total cost of the dining room set be if she must pay off the dining room set within two years after the deferment period?

OpenStudy (anonymous):

@jim_thompson5910 @Hero @Directrix @campbell_st

OpenStudy (campbell_st):

so is the interest simple of compound..?

OpenStudy (anonymous):

must be compounded

OpenStudy (campbell_st):

ok... so the compound interest formula is \[A = P ( 1 + \frac{r}{100})^n\] A = future value, P = principal or initial amount, r = interest rate and n = number of time periods. so I'd assume that n = 2 and r = 23.45 and P = $1730 substitute them into the equation can calculate the actual cost( future value) of the dining room set

OpenStudy (anonymous):

mm.. yah i got it so it would be: 2,754.72 the final answer :)

OpenStudy (anonymous):

thxss man :)

OpenStudy (campbell_st):

I haven't done the calculation but its around $425 more using terms

OpenStudy (anonymous):

yah.. i did all the calculation.. like, u gave me what i was missing of knowledge to do it by myself.. thanks again :)

OpenStudy (campbell_st):

glad to help

OpenStudy (anonymous):

wanna help me with some more? (i can tag you)